# PROPHET: An Inferable Future Forecasting Benchmark with Causal Intervened Likelihood Estimation

### v1

@@ 4.4 Statistics and Properties of PROPHET
We do basic statistics of PROPHET. Assisted by CIL, we also explore key properties of future forecasting task and the benchmark. We currently harvest 99 data in L1 and 53 data in L2. The statistics of news articles we crawled are shown in Table 1. During the construction process, we only discard obviously irrelevant news. Therefore, we did not significantly alter the data distribution of the valid news. News we remain can reflect the real distribution of situations about certain queried events.
We retain 100 top relevant news for each question. The average news tokens are 853.95 leading to context problem if a method longs for simply adding all news in the prompt. We calculate the time span between the oldest and the newest news. The average time span is 16.54 months which is large enough for the method to retrieve similar events in the history for answering.
We conduct in-depth analysis and draw findings of PROPHET assisted by CIL:1) As the resolved date approaches, both high and low CIL  news articles increase. It poses a challenge for models to resist forecasting bias. 2) Two main volume distributions of news articles were identified: one with few articles early on and a sudden surge near the end, and another with a uniform distribution over time. We leave details in the Appendix A.3.
5 Experiments
We first conduct experiments to show the validity of CIL estimation and our benchmark in Section 5.2. Then we evaluate the current retrieval and reasoning baselines on PROPHET in Section 5.3. Lastly, assisted by CIL, we conduct a temporal analysis on PROPHET to provide insights into future forecasting systems in Section 5.4. We use the cases to show the effectiveness of CIL in the Appendix A.6.
5.1 Evaluated Methods
For retrieval methods, we evaluate Naive RAG, APP Halawi et al. (2024), Rankllama Ma et al. (2024), HyDE Gao et al. (2023). For reasoning methods, we include ScrathPAD Halawi et al. (2024), CoT Wei et al. (2022), Long-CoT OpenAI (2024). Details are in the Appendix A.4. Since the news would be long, we pre-summarize each news and all methods use the same summarization in RAG.
Figure 3: Retrieval evaluation.
5.2 Validity of CIL and PROPHET
To validate the estimation of CIL, we conduct branches of experiments. We test numerous methods and LLMs on both L1 and L2 parts of data. The results are shown in Figure 2. To ensure comparability, all methods are on ScratchPAD reasoning prompting. Native RAG and APP are two RAG methods. We also report the differences between w.o. RAG and each RAG method.
As shown, all RAG methods applied to various LLMs perform better than w.o. RAG on L1 while showing little or no improvement on L2. These results strongly suggest that CIL estimation is effective in identifying inferable data. It can measure the supportiveness of news articles. Questions lacking supportive rationales are difficult to accurately forecast. In addition, the results also show CIL estimation is model-agnostic. Although we use GPT-4o to calculate CIL, all models are subjected to these data partitions by CIL. That demonstrates the nature of the intervened causality captured by this robust estimation. Last, we also notice that, in some methods or LLMs, it drops compared to w.o. RAG. It indicates some articles would contribute negatively in prediction. This is consistent with the findings in Section 4.4. Our CIL score is able to measure the negative effects of the news articles.
5.3 Performances on Future Forecasting
In this section, we evaluate current methods in our future forecasting benchmark. We evaluate two branches of methods representing two core abilities of this task, retrieval and reasoning.
5.3.1 Retrieval Performances
We compare between Naive RAG, APP, HyDE, and Rankllama as retrieval evaluation. For all methods, we retrieve 10 news articles and use ScratchPAD reasoning on GPT-4o. We also compare these methods to CIL-high [^5] Note that CIL-high and CIL-low are not actual methods, they are only empirical methods for studying the performance bounds. and CIL-low where we directly use the news articles with the highest and lowest CIL scores. The results are in Figure 3.
CIL-high performs the best while CIL-low is the worst. This further demonstrates the validity of CIL estimation. Among other methods, Rankllama performs the best in Brier Score and improves on CIL score. Rankllama can understand the complicated instructions indicating that it requires deep comprehension of retrieval queries for news. This provides insights that training retrieval methods for complicated query instructions are crucial in such RAG task with hidden rationales.
In total, compared to the CIL-high, all methods still have a significant gap on CIL and Brier Score, indicating that there is still much room for improvement in this retrieval task. It requires delicate approaches that excel in real-world knowledge grounding and comprehension.
Figure 4: Temporal analysis. The horizontal axis represents the entire prediction process.
5.3.2 Reasoning Performances
Reasoning Model $N=5$ $N=10$
CIL-High Naive RAG CIL-High Naive RAG
ScratchPad GPT-4o 17.02 $\pm$ 0.46 21.53 $\pm$ 0.35 16.03 $\pm$ 0.21 21.22 $\pm$ 0.30
GPT-4o-mini 19.37 $\pm$ 0.31 23.66 $\pm$ 0.24 18.37 $\pm$ 0.67 24.03 $\pm$ 0.57
Claude-3-5-sonnet 20.03 $\pm$ 0.17 24.64 $\pm$ 1.16 15.82 $\pm$ 0.53 23.46 $\pm$ 0.85
Gemini-1.5-pro 16.89 $\pm$ 0.35 22.51 $\pm$ 0.19 17.69 $\pm$ 0.54 22.18 $\pm$ 0.39
Qwen2.5-32B 21.38 $\pm$ 1.30 25.10 $\pm$ 0.70 20.74 $\pm$ 1.51 23.89 $\pm$ 0.26
Qwen2.5-7B 26.17 $\pm$ 0.69 30.93 $\pm$ 1.36 24.86 $\pm$ 0.35 26.64 $\pm$ 0.76
CoT GPT-4o 16.70 $\pm$ 1.15 22.04 $\pm$ 0.37 15.60 $\pm$ 0.25 23.75 $\pm$ 0.25
Gemini-1.5-pro 17.68 $\pm$ 0.13 26.45 $\pm$ 2.87 15.57 $\pm$ 1.77 25.34 $\pm$ 1.14
Qwen2.5-32B 17.90 $\pm$ 2.51 22.29 $\pm$ 0.16 15.89 $\pm$ 3.45 26.38 $\pm$ 0.72
Qwen2.5-7B 23.04 $\pm$ 1.87 33.13 $\pm$ 3.60 23.27 $\pm$ 0.42 34.82 $\pm$ 1.33
Long-CoT O1-mini 15.66 $\pm$ 1.14 23.49 $\pm$ 2.94 13.72 $\pm$ 0.38 24.19 $\pm$ 0.65
Table 3: Reasoning evaluation. We report mean and std values on twice runs.
In this section, we evaluate three reasoning methods on PROPHET:ScratchPad, CoT, and Long-CoT. We use various models and test under two retrieval conditions: (1) using news articles with top CIL scores, and (2) using Naive RAG. We also compare retrieval sizes ($N$=5 vs. $N$=10). Results are shown in Table3. Key findings include:
Long-CoT achieves the best results across all methods and models, highlighting its potential for future forecasting tasks. This suggests that event prediction relies heavily on deep, multi-step reasoning based on available information. Specialized post-training in forecasting reasoning is crucial for improving performance.
② Effective information retrieval is fundamental for reasoning. Under Naive RAG, methods show significantly lower performance gains compared to CIL-High. Moreover, models and methods exhibit minimal differences in Naive RAG, while clear distinctions emerge inCIL-High. This underscores the importance of retrieval quality for reasoning. More sophisticated retrieval and reasoning techniques could enhance performance.
ScratchPad outperforms CoT under Naive RAG, but the reverse is true for CIL-High. This finding, not previously reported Halawi et al. (2024), suggests that textttScratchPad constrains the model’s reasoning when useful information is scarce leading to improvements. However, when information is abundant, it may limit the model’s reasoning ability. This insight offers potential for developing advanced reasoning methods.
5.4 Temporal Studies
@@
Future forecasting is a continuous process that begins when the question is posed and ends when the question is answered. The earlier the answer can be predicted, the more valuable it is. We investigate the system’s forecasting at different times. Similar as in Section 4.4, we compute the progress in the whole forecasting. We represent the progress of each news by the percentage of its date in the forecasting. We show the performances of Naive RAG and CIL-High at different times. These experiments are on both L1 part and the whole benchmarks (L1+L2). (L1+L2) is the real-world forecasting scenario. All results are on GPT-4o and ScratchPAD reasoning. The results are in Figure 4.
① We find significant potential in the early-time future forecasting. The CIL-High at 20% progress performs even better than Naive RAG at 100%. It indicates that if we have a sufficiently powerful retrieval method, we can expect to achieve effective predictions at the early stages of event development. This finding applies to both scenarios where evidence is sufficient and where it is insufficient.
② When the forecasting progress precedes, there would be news that is harmful for prediction. We find that during the progress of forecasting, the performances of some methods fluctuate. And the CIL  of the Naive RAG stops increasing at 60%. This is consistent with the conclusions in Section 4.4. It shows a desired prediction system should be aware of negative evidence and can self-correct in the retrieval and reasoning process.
6 Conclusion
@@
Limitations
In this work, we evaluate methods of retrieval and reasoning disentangling. However, entangled methods could further improve future forecasting. We leave it to future work.
Ethics Statement
This dataset is strictly for non-commercial research purposes under the following conditions: 1) Restricted Application Scope: All narrative scenarios contained herein are intended solely for academic exploration of future forecasting methodologies. Any utilization for purposes involving defamation, harassment, malicious targeting, or other unethical practices is expressly prohibited. 2) Prohibited Misinterpretation: Statistical patterns derived from this resource should not be interpreted as deterministic predictions of real-world events. 3) Accountability Framework: The creators explicitly disclaim liability for consequences arising from dataset misuse, including but not limited to algorithmic bias propagation, privacy infringements, or sociotechnical harms caused by improper application.
Appendix A Appendix
A.1 Proof of Proposition
We show the proof of Proposition Eq.(7) below. This proof is mainly based on causal inference theory and our assumptions shown in Figure 2.
Proof.
By the law of total probability,
$P({\mathcal{Y}}=\hat{{\mathcal{Y}}}|do({\mathcal{X}}_{i}=1))\doteq P_{m}({\mathcal{Y}}|{\mathcal{X}}_{i}=1)$ (8)
$=$ $\sum_{n_{1},}\cdots\sum_{m_{1}}\cdots$
$P_{m}({\mathcal{Y}}=\hat{{\mathcal{Y}}}|{\mathcal{X}}_{i}=1,{\mathcal{X}}_{n_{1}},\cdots,{\mathcal{X}}_{m_{1}},\cdots)$
×$\times P_{m}({\mathcal{X}}_{n_{1}},\cdots,{\mathcal{X}}_{m_{1}},\cdots|{\mathcal{X}}_{i}=1)$
$0<\forall n_{j},\mathrm{G}({\mathcal{X}}_{i})-\mathrm{G}({\mathcal{X}}_{n_{j}})\leqq w,$
$\forall m_{j},\mathrm{G}({\mathcal{X}}_{i})-\mathrm{G}({\mathcal{X}}_{m_{j}})>w.$
Since the ${\mathcal{Y}}$ is the latest variable and happened $w$ window later than ${\mathcal{X}}_{i}$, with Assumption 2, we have
$P_{m}({\mathcal{Y}}=\hat{{\mathcal{Y}}}|{\mathcal{X}}_{i}=1,{\mathcal{X}}_{n_{1}},\cdots,{\mathcal{X}}_{m_{1}},\cdots)$ (9)
$=$ $P_{m}({\mathcal{Y}}=\hat{{\mathcal{Y}}}|{\mathcal{X}}_{i}=1,{\mathcal{X}}_{n_{1}},\cdots),$
×$\times P_{m}({\mathcal{X}}_{n_{1}},\cdots,{\mathcal{X}}_{m_{1}},\cdots|{\mathcal{X}}_{i}=1)$
$=$ $P_{m}({\mathcal{X}}_{n_{1}},\cdots|{\mathcal{X}}_{i}=1,{\mathcal{X}}_{m_{1}},\cdots)$
×$\times P({\mathcal{X}}_{m_{1}},\cdots|{\mathcal{X}}_{i}=1)$
$=$ $P_{m}({\mathcal{X}}_{n_{1}},\cdots|{\mathcal{X}}_{i}=1)P({\mathcal{X}}_{m_{1}},\cdots)$
$0<\forall n_{j},\mathrm{G}({\mathcal{X}}_{i})-\mathrm{G}({\mathcal{X}}_{n_{j}})\leqq w,$
$\forall m_{j},\mathrm{G}({\mathcal{X}}_{i})-\mathrm{G}({\mathcal{X}}_{m_{j}})>w.$
Then take Equation (9) into Equation (8), and interchange the order of summation,
$P({\mathcal{Y}}=\hat{{\mathcal{Y}}}|do({\mathcal{X}}_{i}=1))\doteq P_{m}({\mathcal{Y}}|{\mathcal{X}}_{i}=1)$ (10)
$=$ $\sum_{n_{1},}\cdots P_{m}({\mathcal{Y}}=\hat{{\mathcal{Y}}}|{\mathcal{X}}_{i}=1,{\mathcal{X}}_{n_{1}},\cdots)$
×$\times P_{m}({\mathcal{X}}_{n_{1}},\cdots|{\mathcal{X}}_{i}=1)\sum_{m_{1}}\cdots P({\mathcal{X}}_{m_{1}},\cdots)$
$=$ $\sum_{n_{1},}\cdots P_{m}({\mathcal{Y}}=\hat{{\mathcal{Y}}}|{\mathcal{X}}_{i}=1,{\mathcal{X}}_{n_{1}},\cdots)$
×$\times P_{m}({\mathcal{X}}_{n_{1}},\cdots|{\mathcal{X}}_{i}=1)$
$0<\forall n_{j},\mathrm{G}({\mathcal{X}}_{i})-\mathrm{G}({\mathcal{X}}_{n_{j}})\leqq w,$
$\forall m_{j},\mathrm{G}({\mathcal{X}}_{i})-\mathrm{G}({\mathcal{X}}_{m_{j}})>w.$
Under the $do$ operation, ${\mathcal{X}}_{i}$ is independent to ${\mathcal{X}}_{n_{j}},\forall n_{j}$. Owing to Assumptions 1 and 3, the concurrent and later variables don’t influence ${\mathcal{X}}_{i}$. Therefore, the intervened distribution equals to origin distribution.
$P({\mathcal{Y}}=\hat{{\mathcal{Y}}}|do({\mathcal{X}}_{i}=1))\doteq P_{m}({\mathcal{Y}}|{\mathcal{X}}_{i}=1)$ (11)
$=$ $\sum_{n_{1},}\cdots P_{m}({\mathcal{Y}}=\hat{{\mathcal{Y}}}|{\mathcal{X}}_{i}=1,{\mathcal{X}}_{n_{1}},\cdots)P_{m}({\mathcal{X}}_{n_{1}},\cdots)$
$=$ $\sum_{n_{1},}\cdots P({\mathcal{Y}}=\hat{{\mathcal{Y}}}|{\mathcal{X}}_{i}=1,{\mathcal{X}}_{n_{1}},\cdots)P({\mathcal{X}}_{n_{1}},\cdots)$
$\forall n_{j},0<\mathrm{G}({\mathcal{X}}_{i})-\mathrm{G}({\mathcal{X}}_{n_{j}})\leqq w.$
A.2 Construction Details

### v2

@@ 3.5. Statistics and Properties of PROPHET
We perform basic statistics of PROPHET, and, assisted by CIL, we further explore the key properties of the future forecasting task. Currently, we harvest 612 data points in L1 and 43 data points in L2. The statistics of the crawled news articles are shown in Table 1. During the construction process, we only discard obviously irrelevant news, meaning that we did not significantly alter the data distribution of valid news. Thus, the retained articles can reflect the real distribution of situations surrounding queried events.
For each question, we retain 100 top relevant news articles, resulting in an average of 559.51 articles per question. The average token count per article is 1249.88, which poses a challenge for methods that attempt to simply concatenate all news into a single prompt. The average CIL score is 0.013-0.013, indicating that slightly more negative-supportive news exists compared to positive ones.
Beyond these summary statistics, we conduct more in-depth analyses, visualized in Figure 3 and reveal several key insights:
  • Article count distribution. Figure 3(a) shows that most questions have between 400 and 800 associated articles, with a long tail of dense coverage for certain high-profile events. This indicates an uneven information supply across questions.
  • CIL value distribution. As shown in Figure 3(b), the distribution of CIL  values is sharply peaked around zero, indicating that most of the events are non-supportive. It’s hard for methods to mine useful information for precise forecasting.
  • Temporal density of positive/negative articles. Figure 3(c) illustrates article density over the course of event progress (date ranging from the earliest to the latest news) for positive (CIL $\geq 0$) and negative (CIL <0<0) articles. We observe peaks at early stages, implying heightened reporting activity both in the initial phase and just before resolution.
  • Temporal evolution of sentiment. As shown in Figure 3(d), the mean CIL values remain relatively stable over time for both positive and negative groups, with only small oscillations.
Overall, these statistical properties indicate that PROPHET offers both rich contextual diversity and robust inferability, making it a challenging benchmark for forecasting systems that must reason over temporally distributed Web information.
4. Experiments
We first conduct experiments to show the validity of CIL estimation and our benchmark in Section 4.1. Then we evaluate various forecast methods on PROPHET benchmark. We evaluate baselines of both Naive RAG and Agentic RAG.
Naive RAG Naive RAG refers to the straightforward form of retrieval-augmented generation, where the system directly retrieves relevant documents from the knowledge base $\mathbb{X}$} and conditions the generation process on them without iterative reasoning or tool use. Formally, Naive RAG can be represented as:
(11) ${\mathbb{R}}=\mathrm{Retrieve}(Q,B,{\mathbb{X}}),Y=\mathrm{Generate}(Q,B,{\mathbb{R}}),$
where $\mathbb{R}$} denotes the set of retrieved articles, $\mathrm{Retrieve}$ is typically a dense or sparse retriever, and $\mathrm{Generate}$ is an LLM that outputs the probability estimation. In this setting, there is no explicit multi-step reasoning over retrieval results; the model must encode all reasoning based solely on $\mathbb{R}$}. Prompt is in the Appendix prompt (g).
Agentic RAG Agentic RAG extends Naive RAG by introducing an agent that can iteratively retrieve, reason, and act on intermediate conclusions, often following the ReAct paradigm (Yao et al., 2023). In ReAct, the model alternates between reasoning steps $\tau$, action steps $\alpha$ (tool invocation for retrieval), and receiving observation $o$ of the action, enabling multi-step evidence gathering before producing its final probability estimation:
(12) $(Q,\tau_{1},\alpha_{1},o_{1},\tau_{2},\alpha_{2},o_{2},\ldots,\tau_{L},\alpha_{L},o_{L})$
All agents adopt the ReAct framework. Each agent is equipped with the same retrieval tool, prompts, and interaction structure, with the only variation being the underlying LLM. The retrieval tool implements document retrieval: the agent issues a natural language query, which is converted into a dense vector representation via a fixed embedding model, followed by cosine similarity search over the news corpus $\mathbb{X}$} to obtain the most relevant articles. Prompt is in the Appendix prompt (h). The tool is in the Appendix Agentic RAG Tool.
We show the evaluation results in Section 4.3 and 4.4. Then we soly evaluate the reasoning abilities of each model in Section 4.2 and analyze the reasoning performances along the forecasting timeline in Section 4.5. We finally demonstrate cases in Section Case Studies.
4.1. Validity of CIL and PROPHET
To validate the effectiveness of our proposed CIL metrics, we design experiments to test their practical impact. The underlying intuition is that, if CIL is indeed effective, forecasts based on the articles with the highest CIL scores should significantly outperform forecasts obtained by directly answering the question without reference to these articles. To examine this, we conduct hypothesis tests to determine whether there is a statistically significant difference between the two approaches.
We partition the entire news corpus into two disjoint subsets according to the distribution of CIL scores. For each subset we evaluate two strategies: (1) Top CIL: retrieve the top-20 articles ranked by CIL score and feed them into the generator; (2)without RAG: let the generator condition solely on the query.
Every question is repeated 10 times to obtain stable estimates.
Let $X=(x_{i})_{1\times 10n}$, where each $x_{i}$ is the Berrir score of a single response to one question under the with-RAG strategy; analogously, let $Y=(y_{i})_{1\times 10n}$ denote the scores produced without-RAG. since every question is repeated 10 times, the vector contains $10n$ such individual scores. Define the paired differences $Z=(x_{i}-y_{i})_{1\times 10n}$. We conduct one-sample $t$-tests on $Z$ with significance level $\alpha=0.05$ under two one-sided null hypotheses:
$\mathcal{H}_{0}^{\leq}\colon\mu(Z)\leq 0.077\qquad\text{and}\qquad\mathcal{H}_{0}^{\geq}\colon\mu(Z)\geq 0.027.$
Rejecting $\mathcal{H}_{0}^{\leq}$ implies that the average Berrir score of with-RAG significantly exceeds that of without-RAG by at least $0.077$, i.e. the retrieved articles are beneficial for answering the question. Rejecting $\mathcal{H}_{0}^{\geq}$ implies that the improvement is no more than $0.027$, i.e. the articles provide negligible gain.
Subset Hypothesis Decision $p$-value $\bar{d}$
L1 $\mathcal{H}_{0}^{\leq}\colon\mu(Z)\leq 0.077$ Rejected 0.0264 0.0835
$\mathcal{H}_{0}^{\geq}\colon\mu(Z)\geq 0.027$ Not rejected 1.0000
L2 $\mathcal{H}_{0}^{\leq}\colon\mu(Z)\leq 0.077$ Not rejected 1.0000 0.0176
$\mathcal{H}_{0}^{\geq}\colon\mu(Z)\geq 0.027$ Rejected 0.0474
Table 2. Hypothesis-testing results for subsets L1 and L2. $\bar{d}=\bar{x}-\bar{y}$ denotes the mean improvement of Top CIL over without RAG.
The statistical test results reveal that, under the L1 setting, the null hypothesis $\mathcal{H}_{0}^{\leq}$ is rejected ($p=0.0264$), indicating that forecasts constructed from the top-$20$ CIL scored articles achieve a significantly larger improvement (over $7.7$) compared with directly answering the question. In contrast, under the L2 setting, $\mathcal{H}_{0}^{\leq}$ cannot be rejected ($p=1.0000$), suggesting that the improvement over direct answers is not statistically significant. This pattern demonstrates that CIL can successfully identify highly supportive news articles. And the construction of PROPHET is valid based on this metric.
4.2. Reasoning Performances
We select the top-10 CIL articles of each question for prediction, and compare to performances without RAG. The results are shown in Figure 4. From the experimental results, it is clear that the Brier Scores of the top-10 CIL selection are significantly better than those achieved without RAG for all tested models. This further demonstrates the effectiveness of the CIL metric in identifying high-quality articles that are capable of boosting forecasting performance.
Moreover, the strong performance observed in the top-10 CIL setting suggests that the degree of answer leakage within the dataset is minimal. If substantial leakage were present, RAG-assisted prediction would not exhibit such improvements over the baseline, as the retrieved content would simply repeat the ground-truth answers rather than genuinely aiding reasoning. Therefore, these results provide additional evidence that our dataset maintains integrity while allowing meaningful performance gains through retrieval.
4.3. Evaluation On Naive RAG Baselines
We evaluate a set of naive Retrieval-Augmented Generation (RAG) baselines over the constructed PROPHET dataset to establish a performance reference and to examine the practical challenges of the task. For the generator component, we select several representative Large Language Models: Claude-sonnet-4, Doubao-1.5, GPT-4o-mini, DeepSeek-v3, and Gemini-2.5-flash. The retrieval component employs seven popular open-source embedding models: all-MiniLM-L6-v2 (AM), msmarco-distilbert-cos-v5 (MDC), msmarco-MiniLM-L6-cos-v5 (MM), msmarco-distilbert-dot-v5 (MDD), msmarco-bert-base-dot-v5 (MBD), instructor-large (IL), and instructor-base (IB). We evaluate each LLM–embedding pair with retrieval sizes $n=10$ and $n=20$, reporting Brier Scores (lower is better). The results are shown in Table 3. From the results, we make the following takeaways:
Limited capability of Naive RAG. Across all evaluated configurations, introducing naive RAG does not consistently outperform the “w.o. RAG” (no retrieval) baseline, and in many cases even leads to performance degradation. This indicates that simply appending retrieved documents to the LLM input, without any filtering, temporal alignment, or causal reasoning, is insufficient for the PROPHET forecasting task. The retrieved context often contains redundant or irrelevant information, and may conflict with the model’s internal knowledge, which can confuse the generator and inflate the Brier Score. Moreover, the challenges of identifying truly predictive evidence from historical data suggest that naive RAG lacks mechanisms to reason about event timelines, domain-specific causal links, or uncertainty, all of which are essential for accurate forecasting.
Figure 4. Reasoning ability evaluation. The red line stands for random results which is 25.0.
Small differences across current embedding models. Changing among the seven embedding models yields only modest performance variation for a given LLM. This suggests that current general-purpose retrievers are not well suited for temporally and causally grounded forecasting, and that more task-specific embedding models may be needed.
Increasing retrieval size has limited benefit. Increasing from $n=10$ to $n=20$ retrieved documents seldom improves performance and sometimes even degrades it. Merely adding more documents can introduce noise and increase reasoning difficulty for the LLM. High-quality selection of truly relevant evidence is more important than raw retrieval quantity.
Overall, naive RAG configurations underperform on PROPHET, and the results highlight the necessity of causally aware retrieval strategies and reasoning methods tailored to forecasting.
w.o. RAG AM MDC $M$ MDD MBD IL IB
Claude-4-sonnet 18.57 19.00/19.49 19.15/18.72 18.12/18.00 19.15/19.64 18.41/18.87 18.24/19.45 19.69/19.14
Doubao-1-5 22.02 20.73/21.01 21.45/21.00 22.72/21.09 20.53/20.25 21.89/21.23 20.33/21.18 21.85/22.05
GPT-4o-mini 24.28 27.17/27.20 28.32/28.92 28.34/27.89 29.41/28.95 29.08/30.34 27.52/28.06 29.10/28.46
DeepSeek-v3 20.37 21.56/21.48 21.94/21.74 22.22/21.04 21.33/22.60 22.10/22.27 21.60/22.96 23.08/22.55
Gemini-2.5-flash 21.82 21.63/21.48 23.69/22.19 22.58/23.63 21.11/22.68 22.06/20.15 21.54/21.16 21.99/22.06
Table 3. Naive RAG on the PROPHET dataset. Lower is better. Each cell reports the score with n=10n=10/$n=20$ articles.
4.4. Agentic RAG Performances
w.o. RAG AM MDC
BS BS ($\delta$) Step Thought BS ($\delta$) Step Thought
Claude-4-sonnet 18.57 18.65 (+0.08) 3.67 195.65 17.89 (-0.68) 3.58 211.99
Claude-3.5-sonnet 22.22 21.53 (-0.69) 1.05 169.75 21.54 (-0.68) 1.02 202.83
Doubao-1.5 20.20 20.40 (+0.20) 1.02 129.97 24.53 (+4.33) 1.08 166.08
Gemini-25-pr0 21.41 20.51 (-0.90) 0.95 139.64 19.26 (-2.15) 0.95 151.01
GPT-4.1 21.64 21.14 (-0.50) 0.61 91.24 18.44 (-3.20) 0.18 101.94
GPT-4o 22.31 23.33 (+1.02) 1.02 102.64 22.47 (+0.16) 1.04 125.38
Table 4. Performances on Agentic RAG. Lower is better. $\delta$ stands for differences between Agentic RAG and w.o. RAG. Thought is the average thought length of each tool call.
We also conduct experiments on Agentic RAG methods. AM and MDC are different embedding models used in the tool. From Table 4, Agentic RAG achieves consistent gains over the w.o. RAG baseline and outperforms naive RAG (Section 4.3) for several LLMs, with models like GPT-4.1, and Gemini-2.5-pro showing Brier Score reductions exceeding 2.0-2.0. Unlike naive RAG’s simple document concatenation, Agentic RAG allows the LLM to iteratively retrieve, inspect, and integrate evidence, improving temporal and causal relevance while reducing noise. Performance gains are often linked to deeper reasoning, as indicated by higher Step counts and longer Thought lengths. These results highlight Agentic RAG as a promising direction for forecasting, combining adaptive retrieval with structured reasoning to better exploit external knowledge.
4.5. Temporal Analysis
@@
Future forecasting is a continuous process that begins when the question is posed and ends when the question is answered. The earlier the answer can be predicted, the more valuable it is. We investigate the system’s forecasting at different times. Similar to Section 3.5, we compute the progress in the whole forecasting. We represent the progress of each news by the percentage of its date in the forecasting. We run different models based on the top-10 CIL articles in various prediction progress. The results are in Figure 5. As the prediction process progresses, the difficulty of prediction decreases, but early predictions still face great difficulties.
5. Related Work
5.1. Future Forecasting and Benchmarks
Previous research on future forecasting benchmarks has evolved in different paradigms, each addressing different aspects of the task. Early benchmarks, such as MCNC (Granroth-Wilding, 2016), SCT (Mostafazadeh et al., 2017), and CoScript (Yuan et al., 2023), focused on script learning and common sense reasoning in synthetic scenarios. Although these data sets facilitated structured reasoning, they lacked real-world applicability and grounding in factual news. Time series datasets such as GDELT (Leetaru and Schrodt, 2013) and ICEWS (Schrodt et al., 2012) introduced real-world event tracking but did not formalize prediction as a retrieval-augmented reasoning task or ensure answerability. Later works, such as ECARE (Du et al., 2022) and EV2 (Tao et al., 2024), advanced event reasoning has made significant progress in understanding abstract or synthetic scenarios but remains largely confined to settings without real-world grounding, limiting its applicability to practical forecasting or causal inference tasks.
Figure 5. Temporal analysis. Results are on top-10 CIL articles on various forecasting progress.
With the rise of LLMs, recent benchmarks such as Halawi et al. (2024), OpenEP (Guan et al., 2024), and ForecastBench (Karger et al., 2024) shifted the focus to real-world questions and news-based search. However, these datasets suffer from two critical limitations: (1) they lack explicit validation of inferability, allowing questions with insufficient supporting evidence to persist, and (2) they prioritize dynamic data sources over reproducibility, risking inconsistent evaluations due to evolving news archives. PROPHET  addresses these gaps by filtering via the introduced Causal Intervened Likelihood estimation.
5.2. RAG and Benchmarks
Foundational QA Datasets for RAG: Traditional QA datasets, including MMLU (Hendrycks et al., 2021), StrategyQA (Geva et al., 2021), ASQA (Stelmakh et al., 2022), Multi-HopQA (Lin et al., 2020), and 2WikiMultiHopQA (Lin et al., 2020), are adapted to evaluate RAG systems. These datasets, grounded in knowledge bases like Wikipedia, form the basis for RAG evaluation.
Domain-Agnostic: RAGBench (Friel et al., 2024) is a multi-domain benchmark across biomedical, legal, customer support, and finance domains. CRAG (Wang et al., 2024a) provides a factual QA benchmark across five domains, simulating web and knowledge graph search.
Domain-Specific: Domain-specific benchmarks include LegalBench-RAG (Wang et al., 2024b), WeQA (Meyur et al., 2024), PubHealth (Zhang et al., 2023), and MTRAG (Tang and Yang, 2024). These benchmarks address niche applications and improve evaluation precision in domains.
Capability-Oriented: RGB (Liu et al., 2024) evaluates four RAG capabilities: noise robustness, negative rejection, information integration, and counterfactual robustness. TRIAD (Zong et al., 2024) assesses retrieval quality, fidelity, and utility through a three-dimensional framework.
6. Conclusion
@@
Appendix
Case Studies
To evaluate the proposed supportiveness metric in RAG, we consider two representative cases shown in Figure 6. The first question asks whether the total crypto market cap will exceed \$4 trillion on January 20, 2025. Among the retrieved articles, the blue-highlighted news discusses substantial rebounds in Bitcoin and Ether alongside ETF approvals—information directly indicative of a “Yes” outcome. Conversely, the white-highlighted news describes global market declines and cryptocurrency crashes, contradicting the supportive stance. Our metric assigns higher scores to the supportive article, reflecting its stronger evidential link to the answer.
The second question concerns whether Joe Rogan will attend the presidential inauguration. The supportive article details Rogan’s past political activities and his collaboration with Donald Trump during the campaign, which strengthens the likelihood of his attendance. In contrast, the non-supportive article mentions Rogan only in the context of broader shifts in media influence, without providing any concrete statement or event suggesting his presence at the inauguration. Our metric correctly ranks the supportive article higher, aligning with the ground truth answer “Yes”.
Across both cases, the metric effectively distinguishes articles offering direct, answer-relevant evidence from those that are tangential or context-only, thereby enhancing RAG answer accuracy.
Construction Details